6sin^2x - sinx - 2 = 0
Answer by AnlytcPhil(1810)

6sin²x - sinx - 2 = 0 for 0° £ x < 360°

Replace sinx by w, and sin²x by w²

6w² - w - 2 = 0

(2w+1)(3w-2) = 0

2w+1 = 0  ;  3w-2 = 0
  2w = -1      3w = 2
   w = -1/2     w = 2/3

Replace w by sinx

sinx = -1/2        sinx = 2/3

x = 210°±360°k     x = 41.8103149°±360°k
x = 330°±360°k     x = 138.1696851°±360°k
where k = any integer

Edwin
AnlytcPhil@aol.com